OpenStudy (anonymous):

What is the solution to the equation log 2 x – log 2 4 = 2

OpenStudy (anonymous):

log[ 2 ]x – log [2] 4 = 2

OpenStudy (lgbasallote):

log_2 (x/4) = 2 2^2 = x/4 4 = x/4 16 = x

OpenStudy (anonymous):

thankkk youuuuuuu